Assume that the banking system is loaned up and that any open-market purchase by the Fed directly increases reserves in the banks. If the required reserve ratio is 0.2, by how much could the money supply expand if the Fed purchased $2 billion worth of bonds?
1
Expert's answer
2021-08-02T00:48:04-0400

If the required reserve ratio is 0.2, then the money multiplier is:

m = 1/0.2 = 5.

The money supply will expand by:

2×5 = $10 billion.